OpenStudy (anonymous):

Quadratic Equations: A wire of length 200cm is cut into two parts and each part is bent to form a square. If the area of the larger square is 9 times that of the smaller square, find the perimeter of the larger square

OpenStudy (anonymous):

let x cm be the length of the longer wire length of shorter wire is thus (200-x)cm \[(\frac{ x }{ 4 })^{2}= (\frac{ 200-x }{ 4 })^{2}\times 9\] solve for x and you get the length of the longer wire which is in fact the perimeter of the larger square

OpenStudy (anonymous):

do you understand the equation?

OpenStudy (anonymous):

@irkiz I'm having trouble solving the equation

OpenStudy (anonymous):

If the ratio of the areas is 9, then the ratio of the sides is 3. Let x = the length of the smaller square, and 3x = the length of the larger square 4x + 4(3x) = 200 4x + 12x = 200 15x = 200 x = 40/3 3x = 40 The perimeter of the larger square is 160 cm.

OpenStudy (anonymous):

ok @BelleFlower

OpenStudy (anonymous):

Thank you so much! :) @Hares333